OpenStudy (delahruza):

hello, i have an eq: (y^2-x^2-z)*(dx/dt)+(z-y^2-x*y)*(dy/dt)-x*(dz/dt) = 0 how to show that the eqn is integrable?

OpenStudy (anonymous):

what do you mean by integrable, capable of integrating?

OpenStudy (anonymous):

$$\Large f_x \cdot \frac{dx}{dt} + f_y \cdot\frac{dy}{dt}+f_z \cdot\frac{dz}{dt} = 0 $$

OpenStudy (delahruza):

i see now!!! what a blind man i am :( thanks jayzd!!